# Eaman Jahani

### v5

@@ Abstract
As generative AI systems rapidly improve, a key question emerges: How do users keep up—and what happens if they fail to do so. Drawing on theories of dynamic capabilities and IT complements, we examine prompt adaptation—the adjustments users make to their inputs in response to evolving model behavior—as a mechanism that helps determine whether technical advances translate into realized economic value. In a preregistered online experiment with 1,893 participants, who submitted over 18,000 prompts and generated more than 300,000 images, users attempted to replicate a target image in 10 tries using one of three randomly assigned models: DALL-E 2, DALL-E 3, or DALL-E 3 with automated prompt rewriting. We find that users with access to DALL-E 3 achieved higher image similarity than those with DALL-E 2—but only about half of this gain (51%) came from the model itself. The other half (49%) resulted from users adapting their prompts in response to the model’s capabilities. This adaptation emerged across the skill distribution, was driven by trial-and-error, and could not be replicated by automated prompt rewriting, which erased 58% of the performance improvement associated with DALL-E 3. Our findings position prompt adaptation as a dynamic complement to generative AI—and suggest that without it, a substantial share of the economic value created when models advance may go unrealized.
@@ 1 Introduction
Generative AI is being integrated into work practices across the economy (Zhang and Kamel Boulos 2023, Bright et al. 2024), yielding notable productivity gains in tasks as diverse as software development, writing, and material discovery (Brynjolfsson et al. 2023, Dell’Acqua et al. 2023, Noy and Zhang 2023, Peng et al. 2023, Yu 2024, Toner-Rodgers 2024). Recent research points to even greater potential ahead, demonstrating advances in automating core scientific processes (Manning et al. 2024), including tasks as complex as chemical research and proving mathematical theorems (Romera-Paredes et al. 2024, Boiko et al. 2023). The adoption of generative AI is also occurring at an unprecedented pace, with recent research showing that approximately 28% of U.S. workers are already using generative AI in their jobs—a rate that significantly outpaces early adoption of personal computers and internet technology at comparable points in their diffusion (Bright et al. 2024, Bick et al. 2024).
@@
As with many other general-purpose technologies, the effectiveness of generative AI depends not only on the technology itself, but on users’ ability to craft inputs that produce high-quality results. To interact with generative AI systems, users provide written instructions—often referred to as prompts—that guide the model’s behavior. These prompts can range from simple commands (e.g., “write a short story about a robot”) to highly detailed specifications tailored to particular outputs (e.g., a series of paragraphs instructing an AI system to implement a complete piece of software). In this way, prompting serves as a complementary skill—one that, like spreadsheet modeling in the early PC era, can determine the productivity impact of a given tool (Brynjolfsson and Hitt 2000).
@@
Prompting has quickly become an area of active research and practice. Scholars have developed taxonomies of prompt engineering techniques (Oppenlaender 2023), documented recurring patterns in prompt construction (Schulhoff et al. 2024), and examined how developers embed prompts into software systems (Liang et al. 2024). Other studies have explored prompting strategies for specific applications, including image generation (Don-Yehiya et al. 2023, Xie et al. 2023) and clinical documentation (Yao et al. 2024). In parallel, practitioners have built prompt libraries, shared tutorials, and developed tools to support prompt design. These developments signal a growing consensus that prompting plays a meaningful role in extracting value from generative AI systems.
@@
Yet despite this consensus, prompting remains understudied as a dynamic practice. Many prompt libraries and tutorials present effective prompts as reusable artifacts. But prompts that work well with one model version may underperform or break entirely with the next (Liang et al. 2024, Meincke et al. 2025). While recent research increasingly views prompting as an adaptive process, empirical evidence remains limited on how these strategies evolve—both as users refine prompts for a single model and as they adjust to model updates—and on how these changes ultimately affect performance. This raises a broader question for individuals and organizations investing in prompting capabilities: Are prompt strategies transferable across model versions, or must they be continually revised to match changing model behavior?
@@
To begin exploring this question, we identify prompt adaptation [^1] We use the phrase “prompt adaptation” to refer specifically to changes in user prompting behavior that arise in response to evolving model capabilities—distinct from the broader practice of prompt engineering, which includes static best practices, libraries, and templates. as a measurable behavioral mechanism through which user-side inputs evolve alongside technical advances. We conceptualize prompt adaptation as a dynamic complement—that is, a user capability that adapts in response to changes in a technological system and is critical to realizing the full economic value of system improvements. In contrast to static complements (e.g., fixed training, prompt templates), dynamic complements emerge through situated use with rapid feedback, respond to model-level change, and may be enabled or suppressed by system design.
@@
To assess the role of prompt adaptation in shaping realized performance—and to separate its contribution from the direct effects of model improvement—we draw on data from a pre-registered online experiment with $1,893{}$ participants. In the study, participants were asked to replicate a target image using prompts submitted to one of three randomly assigned text-to-image models of varying capability: DALL-E 2, DALL-E 3, or DALL-E 3 with an automated large language model (LLM)-based prompt revision. Each participant submitted at least 10 prompts in an effort to reproduce the image as closely as possible, with a large monetary bonus for top performers. By comparing outcomes across arms—and by conducting a posthoc analysis that re-evaluates prompts on alternative models—we estimate the extent to which users adapted their prompts in response to model improvements and how much this adaptation contributed to overall performance.
We find that participants assigned to DALL-E 3 produce significantly more faithful replications than those assigned to DALL-E 2. Importantly, about half of this improvement comes from participants adapting their prompts to exploit the new model’s capabilities—replaying the old DALL-E 2 prompts on DALL-E 3 yields only about half the total improvement. Furthermore, we find that this prompt adaptation is not limited to advanced “prompt engineers”—participants across the full skill distribution benefit from refining their prompts after seeing how the advanced model responds. Finally, we show that an attempt to automate prompt revision via GPT-4 rewriting does not match the performance achieved by manual human adaptation; in fact, it substantially erodes the gains from DALL-E 3. Together, these findings position prompt adaptation as a key mechanism through which users realize the value of rapidly advancing generative AI systems—and as a concrete example of how digital complements must evolve to keep pace with technological change.
@@
In terms of related literature and additional theory, this experiment builds on work in information systems, emphasizing the importance of dynamic, user-driven complements to digital technologies. Research on IT-enabled dynamic capabilities has shown that the value of new systems depends not only on technical infrastructure but on organizations’ ability to reconfigure routines and user behaviors in response to ongoing change (Bharadwaj 2000, Joshi et al. 2010, Teece et al. 1997). Related work on post-adoptive IT use has demonstrated that users often engage only superficially with new systems and that meaningful performance gains tend to emerge only when users experiment with and refine their interaction strategies over time (Jasperson et al. 2005). Recent research on human-AI collaboration further underscores that interface design and task structure shape the degree to which users can learn from and adapt to model behavior (Fügener et al. 2022). And the concept of co-evolution has been introduced to describe how humans and generative AI systems jointly adapt over time, forming interdependent capabilities that neither could realize alone (Böhm and Schedlberger 2023).
@@
We also engage with work on general-purpose technologies, which has long emphasized that the productivity gains from technical advances depend on the development of new human and organizational complements (Brynjolfsson 1993, Brynjolfsson and Hitt 2000, Brynjolfsson et al. 2021, David 1990). We conceptualize prompt adaptation as one such complement—emerging through trial-and-error, accessible across the skill distribution, and potentially hindered when automation misaligns with user intent. In this way, prompt adaptation shapes how and when technical improvements translate into downstream economic value.
Finally, the contribution of the current paper is four-fold. First, we offer direct causal evidence that prompting is not a fixed input but a dynamic capability that co-evolves with model behavior (Böhm and Schedlberger 2023). Second, we show how even non-expert users adapt their prompting in response to model improvements, extending research on IT-enabled dynamic capabilities (Bharadwaj 2000, Joshi et al. 2010, Teece et al. 1997). Third, we document how performance gains emerge through user-initiated trial-and-error rather than static best practices, advancing work on post-adoptive IT use (Jasperson et al. 2005). Finally, we demonstrate that automation intended to reduce user effort can—if misaligned with user intent—undermine the very adaptations that generate value, contributing to ongoing research on human–AI interaction (Fügener et al. 2022, Yao et al. 2024).
@@
The remainder of the paper is organized as follows. We begin by presenting a simple conceptual framework that characterizes how output quality evolves with improvements in model capacity and with users’ corresponding adjustments in effort. We then introduce an experimental design that closely mirrors this framework and describes the data used in our analysis. Next, we present our empirical findings, including a decomposition of the overall effect into components attributable to model improvements versus prompt adaptation, the impact of automated prompt revision, and heterogeneity across user skill levels. Finally, we compare these empirical results to the model’s predictions and conclude by discussing implications for organizations adopting generative AI.
@@ 2 Conceptual Framework
As generative AI systems evolve, a user’s ability to adapt prompts for improved models can become an important source of realized performance gains. We develop a stylized conceptual framework to formalize how overall output quality depends on both the model’s capacity and the user’s skill and effort in prompt writing. Our goal is not to create a fully normative model. Rather, we wish to clarify the distinction between improvements directly attributable to the model itself versus those arising from user adaptation as motivation for our experimental design. Additionally, the framework naturally implies predictions about the returns to prompting skill and the heterogeneity of such skill.
Although our empirical setting focuses on image replication, this framework generalizes to other tasks where users interact with generative models—such as text generation, code assistance, or molecule design. In each case, we expect performance to reflect both model capacity and users’ efforts to refine their inputs. We present the core findings here; a complete exposition of the framework with formal proofs is provided in Appendix E.1.
@@ 2.1 Notation and Problem Setting
Let $\theta\in(0,1]$ represent the model’s capacity to translate prompts into high-fidelity outputs (for instance, how accurately it captures requested details). Let $s\in(0,1]$ denote a user’s baseline skill in prompt engineering, and let $x\geq 0$ be the effort the user expends on writing and refining prompts. Each unit of $x$ incurs a cost $k\,x$ for some $k>0$, which may reflect the time or cognitive load.
[…]
so that users cannot exceed perfect fidelity (quality of 1), and each additional unit of effort yields diminishing returns. The user’s utility is then
@@
The user chooses their effort such that it maximizes the utility. We assume $\theta s>k$, which implies there is a unique interior optimum $x^{*}(\theta,s)$
@@
Two implications immediately follow from equation 2. Optimal quality is increasing with both model capacity
@@
Furthermore, the relationship between the changes in these parameters supplies us with a basic prediction, which we can evaluate empirically in our experiment.
Proposition 1.
As model capacity $\theta$ improves, the effect of user skill on optimal quality diminishes.
$\frac{\partial^{2}Q^{*}}{\partial s\partial\theta}=-\frac{k}{\theta^{2}s^{2}}<0$
In other words, improvements in model capacity reduce the performance gap between high and low-skilled users.
2.2 Decomposition into Model and Prompt Effects
As model capacity $\theta$ increases, so does $x^{*}$ in equation 1, implying that a more capable model not only delivers better output for a fixed prompt but also encourages the user to devote more effort to composing prompts. This distinction highlights both a direct capacity-based effect of an improved model, a model effect, and an indirect effect driven by users adapting their prompts to the new model’s potential, a prompting effect.
To see this more formally, suppose the model is upgraded from capacity $\theta_{1}$ to $\theta_{2}$. Let the old prompting effort be $x^{*}(\theta_{1},s)$ and the new (adapted) prompting effort be $x^{*}(\theta_{2},s)$. Define the total improvement in quality $\Delta Q$ is
$\Delta Q\;=\;Q\Bigl{(}(\theta_{2},s,x^{*}(\theta_{2},s)\Bigr{)}\;-\;Q\Bigl{(}\theta_{1},s,x^{*}(\theta_{1},s)\Bigr{)}.$
We can decompose $\Delta Q$ into

### v6

@@ Abstract
As generative AI systems rapidly improve, a key question emerges: how do users adapt to these changes, and when does such adaptation matter for realizing performance gains? Drawing on theories of dynamic capabilities and IT complements, we study prompt adaptation—how users adjust their inputs in response to evolving model behavior—using a common experimental design applied to two preregistered tasks with 3,750 total participants who submitted nearly 37,000 prompts. We show that the importance of prompt adaptation depends critically on task structure. In a task with fixed evaluation criteria and an unambiguous goal, user prompt adaptation accounts for roughly half of the performance gains from a model upgrade. In contrast, in an open-ended creative task where the space of acceptable outputs is effectively unbounded and quality is subjective, performance improvements are driven primarily by model capability; prompt adaptation plays a limited role. We further show that automated prompt rewriting cannot generally substitute for human adaptation: when aligned with task objectives, it can modestly improve performance, but when misaligned, it can actively undermine the gains from model improvements. Together, these findings position prompt adaptation as a dynamic complement whose importance depends on task structure and system design, and suggest that without it, a substantial share of the economic value created by advances in generative models may go unrealized.
@@ 1 Introduction
Generative AI is being integrated into work practices across the economy (zhang2023generative, bright2024generative), yielding notable productivity gains in tasks as diverse as software development, writing, and scientific research (brynjolfsson2023generative, dell2023navigating, Noy2023, peng2023impact, Yu2024AIImpact). Recent research points to even greater potential ahead, demonstrating advances in automating core scientific processes (Manning2024Automated), including tasks as complex as chemical research and proving mathematical theorems (RomeraParedes2024Math, Boiko2023AutonomousResearch). The adoption of generative AI is also occurring at an unprecedented pace, with recent research showing that approximately 28% of U.S. workers are already using generative AI in their jobs—a rate that significantly outpaces early adoption of personal computers and internet technology at comparable points in their diffusion (bright2024generative, bick2024rapid).
@@
As with many other general-purpose technologies, the effectiveness of generative AI depends not only on the technology itself, but on users’ ability to craft inputs that produce high-quality results. To interact with generative AI systems, users provide written instructions—or prompts—that guide the model’s behavior. These prompts can range from simple commands (e.g., “write a short story about a robot”) to highly detailed specifications tailored to particular outputs (e.g., a series of paragraphs instructing an AI system to implement a complete piece of software). In this way, prompting serves as a complementary skill—one that, like spreadsheet modeling in the early PC era, can determine the productivity impact of a given tool (brynjolfsson2000beyond).
@@
Prompting has quickly become an area of active research and practice. Scholars have developed taxonomies of prompt engineering techniques (oppenlaender2023), documented recurring patterns in prompt construction (schulhoff2024prompt), and examined how developers embed prompts into software systems (liang2024prompts). Other studies have explored prompting strategies for specific applications, including image generation (donyehiya2023, xie2023) and clinical documentation (yao2024clinicians). In parallel, practitioners have built prompt libraries, shared tutorials, and developed tools to support prompt design. These developments signal a growing consensus that prompting plays a meaningful role in extracting value from generative AI systems.
@@
Yet despite this consensus, prompting remains understudied as a dynamic practice. Many prompt libraries and tutorials present effective prompts as reusable artifacts. But prompts that work well with one model version may underperform or break entirely with the next (liang2024prompts, meincke2025prompting). While recent research increasingly views prompting as an adaptive process, empirical evidence remains limited on how these strategies evolve—both as users refine prompts for a single model and as they adjust to model updates—and on how these changes ultimately affect performance. This raises a broader question for individuals and organizations investing in prompting capabilities: Are prompt strategies transferable across model versions, or must they be continually revised to match changing model behavior?
@@
To begin exploring this question, we identify prompt adaptation [^1] We use the phrase “prompt adaptation” in reference to changes in user prompting behavior in response to evolving model capabilities. This is distinct from the broader practice of prompt engineering, which includes static best practices, libraries, and templates, among other things. as a measurable behavioral mechanism through which user-side inputs evolve alongside technical advances. We conceptualize prompt adaptation as a dynamic complement—that is, a user capability that adapts in response to changes in a technological system and is critical to realizing the full economic value of system improvements. In contrast to static complements (e.g., fixed training, prompt templates), dynamic complements emerge through situated use with rapid feedback, respond to model-level change, and may be enabled or suppressed by system design.
Generative AI systems support a wide variety of task structures. Some tasks are steering-oriented, where users aim to reach a clearly defined goal and success, can be measured against an objective benchmark (e.g., write code to solve a specific problem). Others are creative, where goals are open-ended, evaluation is subjective, and improvement depends more on exploration and curation than on precision (e.g., make a beautiful picture). Recent work helps to establish these regimes: vafa2025s investigate the steerability of models, i.e., how effectively users can reach a specified target, whereas orwig2024language and zhou2024generative describe generative AI–assisted creation as a co-creative process of idea generation, filtering, and aesthetic evaluation. In organizational settings, steering tasks (or bounded tasks, herein) dominate operational workflows such as classification, brand-compliant content generation, or reproducing structured materials where quality can be defined ex ante. Creative tasks (or unbounded tasks, herein), by contrast, underpin innovation activities such as logo or campaign design, product ideation, and concept development, where success depends on novelty, appeal, and fit to evolving goals. Understanding how prompting and prompt adaptation operate across these two regimes is essential for organizations deploying generative AI at scale.
@@
To assess the role of prompt adaptation in shaping outcomes across different types of tasks—and to separate its contribution from the direct effects of model improvement—we conducted two pre-registered online experiments with an aggregate sample of 3,750 participants. The first, a bounded task, asks participants to iteratively prompt one of three randomly assigned text-to-image models—DALL-E 2, DALL-E 3, or DALL-E 3 with automated large-language-model–based prompt rewriting—to reproduce a reference image as accurately as possible. It is explicitly designed to capture steering or target-matching behavior. Each participant submitted at least ten prompts and was eligible for a substantial performance-based bonus, incentivizing careful refinement. The second, an unbounded task, captures creative or open-ended behavior: participants design a logo for a hypothetical organization based on a short textual vignette without a specified target image or objective benchmark. Apart from this difference in goal structure, the two tasks, which were conducted on the same custom interface, were identical from the perspective of participants. By comparing outcomes across treatment arms, conducting post-hoc analyses that re-evaluate prompts on alternative models, and contrasting results across the two experiments, we estimate how users adapted their prompts in response to model improvements and how these adaptations contributed to overall performance across both bounded and unbounded settings.
Across both tasks, we find that participants assigned to DALL-E 3 produced significantly better outputs than those assigned to DALL-E 2, but that the sources of these gains depend on task structure. In the bounded image replication task, about half of the gains came from participants adapting their prompts to exploit the new model’s capabilities—replaying DALL-E 2 prompts on DALL-E 3 yields only about half the total improvement. By contrast, in the open-ended logo generation task, performance gains were driven primarily by improvements in model capability, with prompt adaptation accounting for a much smaller share of the overall effect (roughly 7%, compared to over 90% attributable to the model). Importantly, the prompt adaptation we observe does not appear to be limited to advanced “prompt engineers”: benefits from prompt refinement after interacting with more capable models are observable across the outcome distribution. Finally, to assess whether automation can substitute for user-side prompt adaptation, we examine automated prompt revision via GPT-4 rewriting. We find that such automation can either erode or modestly enhance performance depending on alignment with the end user’s goals: automated rewriting substantially reduced gains in the replication task, but slightly improved performance in the creative setting. Together, these findings position prompt adaptation as a dynamic complement whose importance depends on task structure and system design, shaping how advances in generative AI translate into realized value.
@@
In terms of related literature and additional theory, our research builds on work in information systems, emphasizing the importance of dynamic, user-driven complements to digital technologies. Research on IT-enabled dynamic capabilities has shown that the value of new systems depends not only on technical infrastructure but on organizations’ ability to reconfigure routines and user behaviors in response to ongoing change (bharadwaj2000resource, joshi2010changing, teece1997dynamic). Related work on post-adoptive IT use has demonstrated that users often engage only superficially with new systems and that meaningful performance gains tend to emerge only when users experiment with and refine their interaction strategies over time (jasperson2005comprehensive). Recent research on human-AI collaboration further underscores that interface design and task structure shape the degree to which users can learn from and adapt to model behavior (fugener2022cognitive). And the concept of co-evolution has been introduced to describe how humans and generative AI systems jointly adapt over time, forming interdependent capabilities that neither could realize alone (bohm2023co-evolution).
@@
We also engage with work on general-purpose technologies, which has long emphasized that the productivity gains from technical advances depend on the development of new human and organizational complements (brynjolfsson1993productivity, brynjolfsson2000beyond, brynjolfsson2021productivity, david1990dynamo). We conceptualize prompt adaptation as a dynamic complement that co-evolves with model capability, emerges through use rather than formal training, and whose importance depends on task structure and system design. In this way, prompt adaptation shapes how—and under what conditions—technical improvements translate into downstream economic value.
This paper makes three core contributions. First, we conceptualize prompt adaptation as a dynamic complement to improvements in generative AI models; one through which users actively shape how technical advances translate into realized performance. Leveraging a replay-based analysis enabled by our experimental design, we provide direct empirical evidence for this distinction by separating performance gains attributable to model capability from those arising through user-side adaptation, and show that prompt adaptation can account for a substantial share of realized performance improvements. Second, we identify task structure as a key boundary condition for the importance of prompt adaptation: it is a first-order driver of gains in bounded, steerable tasks with fixed objectives, but plays a much more limited role in open-ended creative tasks where quality is unbounded in nature. Finally, we show that automation intended to simplify prompting is not a neutral substitute for user adaptation. We find that automated prompt rewriting can either undermine or complement performance, depending on its alignment with user goals. Together, these contributions clarify when and how user adaptation amplifies model improvements, positioning prompt adaptation as a dynamic complement whose economic importance depends on task structure and system design.
@@
The remainder of the paper is organized as follows. We begin by presenting a simple conceptual framework that characterizes how output quality evolves with improvements in model capacity and with users’ corresponding adjustments in prompting effort. We then describe our experimental design and the two task settings we study—a bounded, steerable image replication task and an open-ended logo generation task—along with the data and techniques used in our analyses. Next, we present our empirical findings, including a decomposition of the overall effect into components attributable to model improvements versus prompt adaptation, distributional effects across outcome quantiles, and the impact of automated prompt revision. We conclude by synthesizing these results and discussing their implications for organizations adopting generative AI.
@@ 2 Conceptual Framework
We first develop a stylized analytical framework to understand how overall output quality depends jointly on model capacity and on users’ prompting behavior. Our goal is not to propose a fully normative model of user–AI interaction, but rather to distinguish improvements directly attributable to the model itself from those arising through user-side prompt adaptation. This distinction motivates our experimental design and yields testable predictions about how performance gains and their distributional impacts change as models improve.
Although our empirical analysis focuses on two tasks—image replication and logo generation—the same conceptual structure applies broadly to settings in which users interact with generative models, including text generation, code assistance, and scientific research. In all such environments, output quality reflects a combination of model capability, user skill, and prompting effort, and total performance improvements can be decomposed into direct (model-driven) and behavioral (prompting-driven) components. The two tasks we study correspond to different regimes of this framework: image replication is a bounded task with clear ceiling (perfectly replicating the image pixel-for-pixel), whereas logo generation is a more open-ended creative task without an optimal outcome. We develop the core logic for the bounded case in the main text and provide complete derivations for both bounded and unbounded formulations in Appendix A. Nearly all of the main results carry over to the unbounded case, with minor differences arising from curvature differences across specifications. [^2] We note that both the bounded exponential and the unbounded logarithmic forms are illustrative rather than general, and are not intended to be literal representations of the specific tasks we study. They adopt simple structures that allow us to decompose total performance gains into a “model effect” (improvements in capacity holding prompts fixed) and a “prompting effect” (improvements from user adaptation holding capacity fixed). This decomposition is a simplifying abstraction that isolates these mechanisms analytically, though in practice they may interact. The comparative statics of $Q(\theta,s,x)$ also depend on its curvature and on how model capacity and skill interact (e.g., multiplicatively versus additively). We use these tractable forms to convey intuition, not to claim functional-form invariance.
@@ 2.1 Notation and Problem Setting
Let $\theta\in(0,1]$ denote the model’s capacity to translate prompts into high-fidelity outputs, $s\in(0,1]$ denote a user’s baseline skill in prompt engineering, and $x\geq 0$ denote the effort the user expends on writing and refining prompts. Each unit of effort incurs a linear cost $c(x)=kx$ for some $k>0$, reflecting the total cognitive and temporal costs of refining one’s prompts. We assume $k$ is small ($k\!\approx\!0$), consistent with the low marginal cost of iterative prompting and using generative AI models more generally (shahidi2025coasean).
We begin with a bounded quality function,
$Q(\theta,s,x)=1-e^{-\theta sx},$
which is increasing in model capacity, user skill, and prompting effort, but exhibits diminishing marginal returns and has a natural upper bound on performance. We define the user’s utility as
@@
The user chooses their effort to maximize the utility. We assume $\theta s>k$—cost is not prohibitive to effort—which implies there is a unique interior optimum $x^{*}(\theta,s)$:
@@
It follows immediately from Equation 2 that optimal quality is increasing with both model capacity
@@
2.2 Decomposition into Model and Prompting Effects
We now consider an upgrade from a model with capacity $\theta_{1}$ to one with higher capacity $\theta_{2}$. As model capacity $\theta$ increases, Equation 1 implies that optimal effort $x^{*}(\theta,s)$ also rises. Thus, an improved model affects performance through two channels: it directly increases output quality for a fixed prompt, and it indirectly encourages users to invest more effort in refining their prompts. We refer to these as the model effect and the prompting effect, respectively.
Formally, let $x^{*}(\theta_{1},s)$ and $x^{*}(\theta_{2},s)$ denote the user’s optimal prompting effort before and after the upgrade. The total improvement in output quality is
This total gain can be decomposed as